OpenStudy (theslytherinhelper):

At what point(s) on the curve x2 + y2 = 9 is the tangent line vertical?

OpenStudy (theslytherinhelper):

I just clicked that answer and it won't go away, so please ignore it.

OpenStudy (theslytherinhelper):

I got as far as the derivative of the equation is (-x/y)?

OpenStudy (loser66):

yup

OpenStudy (theslytherinhelper):

Great! I don't really know what to do after that, though. Do I have to plug in x values or set the equation to 0 or something?

OpenStudy (loser66):

tangent line is vertical when y =0, replace it to original one to get \(x=\pm 3\)

OpenStudy (theslytherinhelper):

Okay, so, just to make sure, this is what I did (I'm not sure I did it right, but I did get +- 3.): \[x^2+y^2 = 9\] \[\sqrt(y^2) = 9- x^2\] \[y=\sqrt(9-x^2)\] \[0=\sqrt(9-9)\] And then when you plug in 3 or -3, it gives you 0.

OpenStudy (loser66):

It is not wrong but what you did doesn't relate to tangent line of a curve. You must take derivative and take limit of it to see how tangent = infinitive.

OpenStudy (theslytherinhelper):

So, would I just take the derivative \[\frac{ -x }{ y }\], set the y to 0, and then plug in the 3?

OpenStudy (loser66):

Need some verbal argument :)

OpenStudy (loser66):

1) find dy/dx 2) limit dy/dx = 0 iff y =0 3) when y =0, x = +/- 3 4) conclusion.

OpenStudy (theslytherinhelper):

Thank you for your help!